OpenStudy (anonymous):

The pulse rate(P beats per minute) and the speed travelled by an athlete (vm/s) are related by the rule P= 60 + 15V, where 0

OpenStudy (anonymous):

lol i figured it out i thought i was wrong so i didn't bother looking at the answers :)

OpenStudy (anonymous):

dP/dt = 15 *dV/dt ===> dP/dt=15*0.5 = 7.5 ===> dP/dt (at v=6) = 7.5